OpenStudy (anonymous):

1. (1 pt) What is the solution to the equation? 8x + 4 = –20 A. x = –3 B. x = –2 C. x = 2 D. x = 3

OpenStudy (anonymous):

here this is what i did: http://www.cymath.com/answer.php?q=8x%2B4%3D-20

OpenStudy (anonymous):

Thnx are you Muslim

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

thats ok

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

dosent work

OpenStudy (anonymous):

easy. 8x+4=-20 first subtract 4 by both sides. your equation will then look like this--> 8x=-24 then divide both sides by 8. your equation will then look like this--> x+-3 your answer should be -3

OpenStudy (anonymous):

i answered it

OpenStudy (anonymous):

\[1-\frac{ z }{ 5}=9\]

OpenStudy (anonymous):

here is what i got: http://www.cymath.com/answer.php?q=1-z%2F5%3D9

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

2q+18=-5q-3

OpenStudy (anonymous):

here is what i got: http://www.cymath.com/answer.php?q=2q%2B18%3D-5q-3

OpenStudy (anonymous):

ok i guess she is all u need

OpenStudy (anonymous):

do you know any vocab word in math that starts with z?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

any other questions?

OpenStudy (anonymous):

2(2-j)=4(j+10)

OpenStudy (anonymous):

here is what i got: http://www.cymath.com/answer.php?q=2(2-j)%3D4(j%2B10)

OpenStudy (anonymous):

any other ones because im going to have to leave soon

OpenStudy (anonymous):

9 times n =12-3 times n

OpenStudy (anonymous):

http://www.cymath.com/answer.php?q=9*n%3D12-3*n

OpenStudy (anonymous):

u guys are awesome thanks so much!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

I'm starting my history now.

OpenStudy (anonymous):

ok, well good luck, i dont know any of that

OpenStudy (anonymous):

k by